solve the equation in the complex number system:
x squared + 12x + 72 = 0

Answer by RAY100(1637)

x^2 +12x +72 =0
.
use quadratic formula
.
a=1,,,b=12,,,,c=72
.
x= [-(12) +/- sqrt{(12)^2 - 4 (1)(72)}]/ 2(1)
.
x=[-12 +/- sqrt(144 -288}]/2
.
x= [-12 +/- sqrt-144}]/2
.
x= [ -12 +/- 12i ] /2
.
x= -6 +/- 6 i
.
check,(-6+6i)^2 +12(-6+6i) +72 =0
.
36 -72i +36i^2 -72 +72i +72 =0,,,,,ok
.
check for -6-6i will be similar
.